OpenStudy (anonymous):

What is the product in lowest terms? 1/6*5/7 A.6/13 B.5/13 C.1/7 D.5/42

OpenStudy (alexandervonhumboldt2):

1/6 * 5/7=(1/5)/(6*7)=5/42

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

can u help with 4 more @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

What is the product in lowest terms? https://static.k12.com/bank_packages/files/media/mathml_7d8e5b20cc6d2d62d1bcdb499d9d5d988477a9d5_1.gif A.3/40 B.3/10 C.3/7 D.7/13

OpenStudy (alexandervonhumboldt2):

4/5*3/8=(4*3)/(5*8)=12/40=3/10

OpenStudy (anonymous):

What is the product in lowest terms? https://static.k12.com/bank_packages/files/media/mathml_c1f9502dff8e5d04cb0d8acac5bc29b3d45e821d_1.gif A.-20/39 B.-10/39 C.-5/39 D.3/25

OpenStudy (alexandervonhumboldt2):

-5/12*8/13=-(5*8)/(12*13)=-10/39

OpenStudy (anonymous):

What is the product in lowest terms? https://static.k12.com/bank_packages/files/media/mathml_4955160e850b609f2ce50803c7062e65e4b30693_1.gif A.-4/25 B.-3/25 C.-2/25 D.1/5

OpenStudy (alexandervonhumboldt2):

-3/10*8/15=-(3*8)/(10*15)=-24/150=-4/25

OpenStudy (anonymous):

Serena bought 3/4 pound of cheese. She used 2/3 of the cheese for sandwiches. How much cheese did Serena use for sandwiches? A. 1/4 lb B. 1/3 lb C. 1/2 lb D. 5/7 lb

OpenStudy (alexandervonhumboldt2):

3/4*2/3=6/12=1/2